OpenStudy (anonymous):

anyone? =( If A (with arrow above)= (6.00i-8.00j) units, B (with arrow above)= (-8.00i+ 3.00j) units, and C (with arrow above) = (26.0i+ 19.0j) units, determine a and b such that aA(with arrow above+ bB(with arrow above) + C(with arrow above) = 0. a= b=

OpenStudy (anonymous):

OK - here goes: a(6i-8j)+b(-8i+3j)+(26i+19j)=0 (6a-8b+26) * i + (-8a+3b+19) * j =0 Of course, 0*i + 0*j = 0 So, equating the coefficients of i to 0 and the coefficients of j to 0 we will get a system of two equations with two variables: 6a-8b+26=0 (multiplier of i equals 0) -8a+3b+19=0 (multiplier of j equals 0) Solving the above equations: --> a=5, b=7